# Training towards significance with the decorrelated event classifier transformer neural network

### v2

@@ Abstract
Experimental particle physics uses machine learning for many of tasks, where one application is to classify signal and background events. The classification can be used to bin an analysis region to enhance the expected significance for a mass resonance search. In natural language processing, one of the leading neural network architectures is the transformer. In this work, an event classifier transformer is proposed to bin an analysis region, in which the network is trained with special techniques. The techniques developed here can enhance the significance and reduce the correlation between the network’s output and the reconstructed mass. It is found that this trained network can perform better than boosted decision trees and feed-forward networks.
@@ I Introduction
Experimental particle physics is often performed by colliding particles and observing their interaction with detectors. Particles generated from the collisions are investigated by reconstructing the detector data. A common method used in particle searches is to search for a resonance in the reconstructed mass distribution. An example of the resonance is shown in Fig. 1. Signal events create a peak that can be seen above background events. To estimate the number signal and background events, the mass distribution is fitted with a peaking signal and smooth background function. The fitted mass region is set to be larger than the signal peak width to estimate the background more precisely. This method is referred to as “bump hunting” and was used to search for the Higgs boson [1, 2]. The sensitivity of the method can be quantified in terms of significance [3, 4]. A larger significance corresponds to a smaller probability that the bump was created by random statistical fluctuations of the background.
Figure 1: (Left) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay, where a bump can be seen due the presence of the Higgs boson particle. The Higgs boson cross section was scaled up by 100 to make the bump visible. (Right) Mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay with the nominal Higgs boson cross section, where the bump can not be seen due to the background.
To increase the expected significance of the bump hunting method, requirements are applied to create a search region to suppress the background, while preserving the signal. The sensitivity of the analysis can be further enhanced by “binning”, where the search region is divided into multiple bins according to other variables. However, if the binning affects the reconstructed mass distribution of the background, estimating the number of signal events can become difficult. An extreme case would be where the background peaks similarly to the signal. Therefore, a desirable characteristic of binning in the other variables is where it does not affect the shape of background distribution. Binning can be performed using distinguishing features of the signal [2] or by using machine learning techniques [1, 5, 6] that classify the signal and background events. A common machine learning technique used for this task is the boosted decision tree (BDT).
@@ II Event classifier transformer neural network
Transformer neural networks [7] are the leading neural network architecture in the natural language processing field. The architecture has been applied to particle physics in a network called Particle Transformer [8]. The network identifies the type of particle that produced a jet, which is a cluster of hadrons whose momentum lie within a cone. In this work, a new transformer-based neural network called event classifier transformer is proposed that classifies signal and background events to bin a bump-hunt analysis.
The architecture of the event classifier transformer is shown in Fig. 2. The inputs are the features of the event, such as the momenta or angles of particles. To be able to use the transformer architecture, each input or each set of inputs is “embedded” into a token, which is a set of $N_{\text{repr.}}$ numbers, using separate feed-forward neural networks. In contrast to the transformer, there is no positional encoding because event features do not have sequential dependence. The tokens are normalized using layer normalization [9] and passed to a multi-headed attention layer (MHA) [7]. The output is summed to employ a residual connection [10] and normalized with layer normalization. A position-wise feed-forward neural network [7] is applied with residual connection and layer normalization, where the same position-wise network is applied to each token. The output is a group of tokens, where each token is a contextual representation of the input feature. Following a simplified version of the CaiT [11] approach and the Particle Transformer, a trainable class token is passed as a query to an MHA and the contextualized tokens concatenated with the class token is passed as a key and value. The output employs a residual connection and layer normalization and is passed to a linear layer. The linear layer output is a single value that is passed through a sigmoid to represent the probability that the event is from the signal process. To prevent overtraining, dropout layers [12] are used in the scores of the MHA and before the residual connection during training. Implementation of the network is publicly available at [13].
Figure 2: Architecture of the event classifier transformer. FFN refers to feed-forward neural network. Normalize refers to layer normalization. Add refers to the implementation of the residual connection. Concat refers to a layer concatenating tokens. Linear refers to linear layer.
@@ III Training techniques for enhancing significance
Special training techniques are developed to apply to the event classifier transformer and other neural networks, to increase the expected significance and reduce the correlation for a bump-hunt analysis. The following new training techniques are investigated:
  • Specialized loss function with mass decorrelation
  • Data scope training
  • Significance based model selection
These techniques are described in turn below, where implementation is publicly available at [13].
@@ III.1 Specialized loss function with mass decorrelation
In a binary classification task, where the input $x$ is provided to predict the label $y$, a neural network $f(x)$ can be trained to output a prediction $\hat{y}=f(x)$. In this work, signal is defined to be $y=1$ and background is defined to be $y=0$. The network is trained with a loss function that is used to minimize the difference between the network’s output $\hat{y}$ with the label $y$.
@@
In this work, to increase the penalty when $\hat{y}$ and $y$ are different, while keeping the property that the minimum is achieved at $\hat{y}=y$, an alternative loss inspired from BCE called extreme loss $E(\hat{y},y)$ is proposed,
@@
For a given value of $\hat{y}-y$, the extreme loss penalizes the neural network more than the BCE loss, as can be seen in Fig. 3. This extreme loss heavily suppresses the background that have network predictions close to 1 and also the signal that have network predictions close to 0. Because the expected significance when binning a bump-hunt analysis is sensitive to backgrounds that have high network predictions, this loss can help increase the significance.
Figure 3: BCE loss vs extreme loss, when the label is $y=0$. Extreme loss penalizes the neural network more than BCE loss for network prediction’s that are close to 1.
To decorrelate the neural network output with the reconstructed mass, Distance Correlation (DisCo) regularization is used [15]. DisCo measures the dependence between $\hat{y}$ and mass, where the value of DisCo is zero if and only if $\hat{y}$ and mass are independent. DisCo is multiplied by a factor $\lambda$ and added to the classifier loss function,
@@ III.2 Data scope training
To increase the significance and decorrelate the neural network output with the reconstructed mass for the background, the loss terms are calculated with different data scopes during training,
@@
The classifier loss, such as BCE loss, is calculated in the scope of signal and background events that fall in the narrow mass window, where the majority of the signal lies. This narrow scope decorrelates the network’s output with the mass and can improve the significance. The DisCo term is calculated in the scope of background events that fall in the wide mass window, where the mass window is used in estimating the amount of background in the bump-hunt method.
III.3 Significance based model selection
When training networks, the best trained network among the training epochs can be chosen based on an evaluation metric. The loss is often used as the evaluation metric, but the network that has the minimum loss can be different from the network that has the best significance [3, 16]. In this work, the expected significance is used as a metric to select the best network among the epochs.
The significance is calculated as followings,
@@ III.3 Significance based model selection
where $N_{S}$ is the number of signal events and $N_{B}$ is the number of background events within a mass window containing the 5 to 95 percentile of the signal. Note that the mass window width can change depending on the bin. The combination of significance [17] over the bins is calculated with
@@ IV Example analysis
The following sections describe the data set, input features for the machine learning techniques, machine learning techniques, evaluation metrics, experiment, and results.
@@ IV.1 Data set
  • Trigger threshold requirements: $p_{T}^{\text{leading }\ell}>25\text{ GeV}$, $p_{T}^{\text{subleading }\ell}>15\ \text{GeV}$
  • Background suppressions requirements: $p_{T}^{\gamma}/m_{\ell\ell\gamma}>15/110$, minimum $\Delta R\left(\ell^{\pm},\gamma\right)>0.4$, $m_{\ell\ell}>50\ \text{GeV}$, $m_{\ell\ell\gamma}+m_{\ell\ell}>185\ \text{GeV}$
  • Mass window requirement: $100<m_{\ell\ell\gamma}<180\ \text{GeV}$
After applying these requirements, there are 9 million events for the signal and 5 million events for the background. The large sample helps to reduce overtraining, where the significance metric is sensitive to the sample region that has high classifier scores and low number of background events.
IV.2 Input features for machine learning techniques
The following features are used as inputs for the machine learning techniques that bin the analysis:
@@ IV.2 Input features for machine learning techniques
  • $p_{Tt}^{\ell\ell\gamma}$: Projection of the reconstructed Higgs boson $p_{T}$ to the di-lepton thrust axis [29].
@@
Many of the features are correlated with the reconstructed Higgs mass. Therefore depending on which features are included in the machine learning technique and which features the machine learning technique prioritizes, the output of machine learning classifier can also be correlated with the mass. Especially when including the momenta of the photon and lepton in certain machine learning techniques, such as BDTs, to bin the search region, the background mass distribution in certain bins tend to peak close to the Higgs boson mass, which can be seen in Fig. 4. This behavior introduces difficulties in estimating the number of signal events, so that these features are typically excluded for the inputs of the machine learning technique. However these features can be used as inputs for neural networks that are trained to be decorrelated with the reconstructed mass.
Figure 4: (Above) Reconstructed $m_{\ell\ell\gamma}$ background distributions, where each histogram is a bin in the XGBoost output distribution with equal number of signal events. Lower signal percentile (sig. p) values correspond to higher output values. $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ are not used in the training. (Below) Reconstructed $m_{\ell\ell\gamma}$ background distributions, when training includes $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ inputs for XGBoost. For certain bins, the backgrounds peaks close to the Higgs boson mass of 125 GeV, which introduces difficulties in estimating the number of signal events. Correlation represents the magnitude of difference in the shapes between the machine learning bins.
IV.3 Machine learning techniques
The following machine learning techniques are compared:
@@ IV.3 Machine learning techniques
  • XGBoost [33] with 100 number of boosting rounds with a maximum depth of 3. The BDT was optimized by modifying the number of boosting rounds, where BDTs with more than 100 round have higher overtraining, while other performance metrics are similar.
@@
  • Feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features, a hidden layer with $4N$ nodes with a $\tanh$ activation function, and an output layer with 1 node with a sigmoid activation function. The network is implemented with PyTorch [34].
@@
  • Deep feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features and three hidden layers with $3N$, $9N$, and $3N$ nodes with a $\tanh$ activation function. The last hidden layer is connected to a dropout layer with a rate of 0.1. The output layer has a single node with a sigmoid activation function. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout had significant overtraining and poorer performance, while the network with a dropout rate of 0.2 had similar performance with the dropout 0.1 network.
@@
  • Event classifier transformer with $N_{\text{repr.}}=16$. Each event feature is embedded by a separate feed-forward network that has an input layer with one node, a hidden layer with 4 nodes with a GELU activation function, and a output layer with 16 nodes. The MHA has 4 heads and the position-wise feed-forward network has an input layer with 16 nodes, a hidden layer with 64 nodes with a GELU activation function, and a output layer with 16 nodes. The dropout layer has a rate of 0.1. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout showed overtraining and poorer performance, while the network with a dropout rate of 0.2 had slighter worse performance than the dropout 0.1 network.
@@
When training the machine learning techniques, the signal and background samples are weighted to have the same number of events. All input features are normalized for neural networks.
@@
For BDTs and networks that are trained without DisCo loss, 12 input features are used: $\eta_{\gamma}$, $\eta_{\ell}^{\text{leading }\ell}$, $\eta_{\ell}^{\text{subleading }\ell}$, minimum $\Delta R\left(\ell^{\pm},\gamma\right)$, maximum $\Delta R\left(\ell^{\pm},\gamma\right)$, flavor of $\ell$, $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$, $p_{Tt}^{\ell\ell\gamma}$, $\sigma_{m}^{\ell\ell\gamma}$, $\cos\Theta$, $\cos\theta$, $\phi$. These input features have minimal correlation with the reconstructed $m_{\ell\ell\gamma}$. The training is performed with events that have a reconstructed Higgs boson mass range between 120 and 130 GeV.
@@
When networks are trained with DisCo loss, 3 additional features are used: $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$. The data scope training method described in section III.2 is used. The narrow mass window is from 120 to 130 GeV and the wide mass window is from 100 to 180 GeV.
Different combinations of loss functions are explored,
@@
    Extreme + DisCo loss
To increase the stability of the calculation, the extreme loss implementation clamps the neural network output $\hat{y}$ to a range from 0.001 to 0.999. The $\lambda$ DisCo factor is 10 for the BCE + DisCo loss and 50 for extreme + DisCo loss to have similar background correlations among networks trained with DisCo.
Each neural network is trained for 1200 training epochs with the Adam optimization algorithm [35] with a learning rate of $10^{-3}$ and a batch size of 8192. The best model between the epochs is selected by finding the model with highest significance on the validation data set. To reduce the amount of time in training, the model is evaluated on the validation data set in 5 training epochs intervals for the first 50 epochs and then 10 training epoch intervals for the remaining training epochs. This method allows the best model search to be sensitive to the early epochs, where the evaluation metric can change dynamically.
IV.4 Machine learning technique evaluation metrics
The machine learning techniques are evaluated with the following metrics:
  • Expected significance
  • Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve
  • Correlation of $m_{\ell\ell\gamma}$ with the machine learning technique output
  • Epoch of the best trained model
The significance is calculated using the method described in section III.3, where the sample is divided into 8 machine learning technique bins with equal number of signal events in each bin.

### v3

# Training toward significance with the decorrelated event classifier transformer neural network

@@ Abstract
Experimental particle physics uses machine learning for many tasks, where one application is to classify signal and background events. This classification can be used to bin an analysis region to enhance the expected significance for a mass resonance search. In natural language processing, one of the leading neural network architectures is the transformer. In this work, an event classifier transformer is proposed to bin an analysis region, in which the network is trained with special techniques. The techniques developed here can enhance the significance and reduce the correlation between the network’s output and the reconstructed mass. It is found that this trained network can perform better than boosted decision trees and feed-forward networks.
@@ I Introduction
Experimental particle physics is often performed by colliding particles and observing their interaction with detectors. Particles generated from the collisions are investigated by reconstructing the detector data. A common method used in particle searches is to search for a resonance in the reconstructed mass distribution. An example of this resonance is shown in Fig. 1. Signal events create a peak that can be seen above background events. To estimate the number of signal and background events, the mass distribution is fitted with a peaking signal and a smooth background function. The fitted mass region is set to be larger than the signal peak width to estimate the background more precisely. This method is referred to as “bump hunting” and was used to search for the Higgs boson [1, 2]. The sensitivity of the method can be quantified in terms of significance [3, 4]. A larger significance corresponds to a smaller probability that the bump was created by random statistical fluctuations of the background.
Figure 1: Left: mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay, where a bump can be seen due to the presence of the Higgs boson particle. The Higgs boson cross section was scaled up by 100 to make the bump visible. Right: mass of reconstructed Higgs boson candidates from the $H\rightarrow Z\left(\ell^{+}\ell^{-}\right)\gamma$ decay with the nominal Higgs boson cross section, where the bump cannot be seen due to the background.
To increase the expected significance of the bump hunting method, requirements are applied to create a search region to suppress the background, while preserving the signal. The sensitivity of the analysis can be further enhanced by “binning”, where the search region is divided into multiple bins according to other variables. However, if the binning affects the reconstructed mass distribution of the background, estimating the number of signal events can become difficult. An extreme case would be one where the background peaks similarly to the signal. Therefore, a desirable characteristic of binning in the other variables is that it does not affect the shape of the background distribution. Binning can be performed using distinguishing features of the signal [2], or by using machine-learning techniques [1, 5, 6] that classify the signal and background events. A common machine-learning technique used for this task is the boosted decision tree (BDT).
@@ II Event classifier transformer neural network
Transformer neural networks [7] are the leading neural network architecture in the natural language processing field. The architecture has been applied to particle physics in a network called Particle Transformer [8]. The network identifies the type of particle that produced a jet, which is a cluster of hadrons whose momenta lie within a cone. In this work, a new transformer-based neural network called an event classifier transformer is proposed that classifies signal and background events to bin a bump hunt analysis.
The architecture of the event classifier transformer is shown in Fig. 2. The inputs are the features of the event, such as the momenta or angles of particles. To be able to use the transformer architecture, each input or each set of inputs is “embedded” into a token, which is a set of $N_{\text{repr.}}$ numbers, using separate feed-forward neural networks. In contrast to the transformer, there is no positional encoding because event features do not have sequential dependence. The tokens are normalized using layer normalization [9] and passed to a multiheaded attention layer (MHA) [7]. The output is summed to employ a residual connection [10] and normalized with layer normalization. A positionwise feed-forward neural network [7] is applied with residual connection and layer normalization, where the same positionwise network is applied to each token. The output is a group of tokens, where each token is a contextual representation of the input feature. Following a simplified version of the CaiT [11] approach and the Particle Transformer, a trainable class token is passed as a query to a MHA and the contextualized tokens concatenated with the class token are passed as a key and value. The output employs a residual connection and layer normalization and is passed to a linear layer. The linear layer output is a single value that is passed through a sigmoid to represent the probability that the event is from the signal process. To prevent overtraining, dropout layers [12] are used in the scores of the MHA and before the residual connection during training. Implementation of the network is publicly available at Ref. [13].
Figure 2: Architecture of the event classifier transformer. FFN refers to a feed-forward neural network. Normalize refers to layer normalization. Add refers to the implementation of the residual connection. Concat refers to a layer concatenating tokens. Linear refers to a linear layer.
@@ III Training techniques for enhancing significance
Special training techniques are developed to apply to the event classifier transformer and other neural networks, to increase the expected significance and reduce the correlation for a bump hunt analysis. The following new training techniques are investigated:
  • Specialized loss function with mass decorrelation.
  • Data scope training.
  • Significance based model selection.
These techniques are described in turn below, and implementation is publicly available at Ref. [13].
@@ III.1 Specialized loss function with mass decorrelation
In a binary classification task, where the input $x$ is provided to predict the label $y$, a neural network $f(x)$ can be trained to output a prediction $\hat{y}=f(x)$. In this work, signal is defined to be $y=1$ and background is defined to be $y=0$. The network is trained with a loss function that is used to minimize the difference between the network’s output $\hat{y}$ and label $y$.
@@
In this work, to increase the penalty when $\hat{y}$ and $y$ are different, while keeping the property that the minimum is achieved at $\hat{y}=y$, an alternative loss inspired from BCE called extreme loss, $E(\hat{y},y)$, is proposed:
@@
For a given value of $\hat{y}-y$, the extreme loss penalizes the neural network more than the BCE loss, as can be seen in Fig. 3. This extreme loss heavily suppresses backgrounds that have network predictions close to 1 and also signals that have network predictions close to 0. Because the expected significance when binning a bump hunt analysis is sensitive to backgrounds that have high network predictions, this loss can help to increase the significance.
Figure 3: BCE loss vs extreme loss, when the label is $y=0$. Extreme loss penalizes the neural network more than BCE loss for network predictions that are close to 1.
To decorrelate the neural network output with the reconstructed mass, Distance Correlation (DisCo) regularization is used [15]. DisCo measures the dependence between $\hat{y}$ and mass, where the value of DisCo is zero if and only if $\hat{y}$ and mass are independent. DisCo is multiplied by a factor $\lambda$ and added to the classifier loss function:
@@ III.2 Data scope training
To increase the significance and decorrelate the neural network output with the reconstructed mass for the background, the loss terms are calculated with different data scopes during training:
@@
The classifier loss, such as BCE loss, is calculated in the scope of signal and background events that fall in the narrow mass window, where the majority of the signal lies. This narrow scope decorrelates the network’s output with the mass and can improve the significance. The DisCo term is calculated in the scope of background events that fall in the wide mass window, where the mass window is used in estimating the amount of background in the bump hunt method.
III.3 Significance-based model selection
When training networks, the best-trained network among the training epochs can be chosen based on an evaluation metric. The loss is often used as the evaluation metric, but the network that has the minimum loss can be different from the network that has the best significance [3, 16]. In this work, the expected significance is used as a metric to select the best network among the epochs.
The significance is calculated with the following steps:
@@ III.3 Significance-based model selection
where $N_{S}$ is the number of signal events and $N_{B}$ is the number of background events within a mass window containing the 5th to 95th percentile of the signal. Note that the mass window width can change depending on the bin. The combination of significance [17] over the bins is calculated with
@@ IV Example analysis
The following sections describe the dataset, input features for the machine-learning techniques, the machine-learning techniques themselves, evaluation metrics, experiment, and results.
@@ IV.1 Dataset
  • Trigger threshold requirements: $p_{T}^{\text{leading }\ell}>25\text{ GeV}$, $p_{T}^{\text{subleading }\ell}>15\ \text{GeV}$.
  • Background suppression requirements: $p_{T}^{\gamma}/m_{\ell\ell\gamma}>15/110$, minimum $\Delta R\left(\ell^{\pm},\gamma\right)>0.4$, $m_{\ell\ell}>50\ \text{GeV}$, $m_{\ell\ell\gamma}+m_{\ell\ell}>185\ \text{GeV}$.
  • Mass window requirement: $100<m_{\ell\ell\gamma}<180\ \text{GeV}$.
After applying these requirements, there are 9 million events for the signal and 5 million events for the background. The large sample helps to reduce overtraining, where the significance metric is sensitive to the sample region that has high classifier scores and a low number of background events.
IV.2 Input features for machine-learning techniques
The following features are used as inputs for the machine-learning techniques that bin the analysis:
@@ IV.2 Input features for machine-learning techniques
  • $p_{Tt}^{\ell\ell\gamma}$: Projection of the reconstructed Higgs boson $p_{T}$ to the dilepton thrust axis [29].
@@
Many of the features are correlated with the reconstructed Higgs mass. Therefore depending on which features are included in the machine-learning technique and which features the machine-learning technique prioritizes, the output of machine-learning classifier can also be correlated with the mass. Especially when including the momenta of the photon and lepton in certain machine-learning techniques, such as BDTs, to bin the search region, the background mass distribution in certain bins tends to peak close to the Higgs boson mass, which can be seen in Fig. 4. This behavior introduces difficulties in estimating the number of signal events, so that these features are typically excluded for the inputs of the machine-learning technique. However, these features can be used as inputs for neural networks that are trained to be decorrelated with the reconstructed mass.
Figure 4: Top: reconstructed $m_{\ell\ell\gamma}$ background distributions, where each histogram is a bin in the XGBoost output distribution with an equal number of signal events. Lower signal percentile (sig. p) values correspond to higher output values. $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, $p_{T}^{\text{subleading }\ell}$ are not used in the training. Bottom: reconstructed $m_{\ell\ell\gamma}$ background distributions, when training includes $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, and $p_{T}^{\text{subleading }\ell}$ inputs for XGBoost. For certain bins, the background peaks close to the Higgs boson mass of 125 GeV, which introduces difficulties in estimating the number of signal events. Correlation represents the magnitude of difference in the shapes between the machine-learning bins.
IV.3 Machine-learning techniques
The following machine-learning techniques are compared:
@@ IV.3 Machine-learning techniques
  • XGBoost [33] with 100 boosting rounds with a maximum depth of 3. The BDT was optimized by modifying the number of boosting rounds, where BDTs with more than 100 rounds have higher overtraining, while other performance metrics are similar.
@@
  • Feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features, a hidden layer with $4N$ nodes with a $\tanh$ activation function, and an output layer with one node with a sigmoid activation function. The network is implemented with PyTorch [34].
@@
  • Deep feed-forward neural network that has an input layer with $N$ nodes corresponding to the number of input features and three hidden layers with $3N$, $9N$, and $3N$ nodes with a $\tanh$ activation function. The last hidden layer is connected to a dropout layer with a rate of 0.1. The output layer has a single node with a sigmoid activation function. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout had significant overtraining and poorer performance, while the network with a dropout rate of 0.2 had similar performance to the dropout 0.1 network.
@@
  • Event classifier transformer with $N_{\text{repr.}}=16$. Each event feature is embedded by a separate feed-forward network that has an input layer with one node, a hidden layer with 4 nodes with a GELU activation function, and an output layer with 16 nodes. The MHA has 4 heads, and the positionwise feed-forward network has an input layer with 16 nodes, a hidden layer with 64 nodes with a GELU activation function, and an output layer with 16 nodes. The dropout layer has a rate of 0.1. The network is implemented with PyTorch. The network was optimized by modifying the dropout rate, where the network with no dropout showed overtraining and poorer performance, while the network with a dropout rate of 0.2 had slightly worse performance than the dropout 0.1 network.
@@
When training the machine-learning techniques, the signal and background samples are weighted to have the same number of events. All input features are normalized for neural networks.
@@
For BDTs and networks that are trained without DisCo loss, 12 input features are used: $\eta_{\gamma}$, $\eta_{\ell}^{\text{leading }\ell}$, $\eta_{\ell}^{\text{subleading }\ell}$, minimum $\Delta R\left(\ell^{\pm},\gamma\right)$, maximum $\Delta R\left(\ell^{\pm},\gamma\right)$, flavor of $\ell$, $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$, $p_{Tt}^{\ell\ell\gamma}$, $\sigma_{m}^{\ell\ell\gamma}$, $\cos\Theta$, $\cos\theta$, and $\phi$. These input features have minimal correlation with the reconstructed $m_{\ell\ell\gamma}$. The training is performed with events that have a reconstructed Higgs boson mass range between 120 and 130 GeV.
@@
When networks are trained with DisCo loss, three additional features are used: $p_{T}^{\gamma}/m_{\ell\ell\gamma}$, $p_{T}^{\text{leading }\ell}$, and $p_{T}^{\text{subleading }\ell}$. The data scope training method described in Sec. III.2 is used. The narrow mass window is from 120 to 130 GeV, and the wide mass window is from 100 to 180 GeV.
Different combinations of loss functions are explored:
@@
    Extreme + DisCo loss.
To increase the stability of the calculation, the extreme loss implementation clamps the neural network output $\hat{y}$ to a range from 0.001 to 0.999. The $\lambda$ DisCo factor is 10 for the BCE + DisCo loss and 50 for the extreme + DisCo loss to have similar background correlations among networks trained with DisCo.
Each neural network is trained for 1200 training epochs with the Adam optimization algorithm [35] with a learning rate of $10^{-3}$ and a batch size of 8192. The best model between the epochs is selected by finding the model with highest significance on the validation dataset. To reduce the amount of time in training, the model is evaluated on the validation dataset in five training epochs intervals for the first 50 epochs, and then ten training epoch intervals for the remaining training epochs. This method allows the best model search to be sensitive to the early epochs, where the evaluation metric can change dynamically.
IV.4 Machine-learning technique evaluation metrics
The machine-learning techniques are evaluated with the following metrics:
  • Expected significance.
  • Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve.
  • Correlation of $m_{\ell\ell\gamma}$ with the machine-learning technique output.
  • Epoch of the best-trained model
The significance is calculated using the method described in Sec. III.3, where the sample is divided into eight machine learning technique bins with an equal number of signal events in each bin.